Tom is driving a car with a mass of 1100kg. He takes it to a racetrack to see how fast he can go around the corner. If there is a banked Curve of 25 degrees and the coefficient of friction between the car and the road is 0.6, what is the fastest Tom can travel?


1
Expert's answer
2021-05-31T17:17:18-0400

"F_1=F_2,"

"F_1=\frac{mv^2}{r},"

"F_2=\mu N=\frac{\mu m g}{\cos\alpha},"

"v=\sqrt{\frac{\mu gr}{\cos \alpha}}=\sqrt{6.5r}~\frac ms."